# Building reliable sim driving agents by scaling self-play

### v1

@@ Abstract
Simulation agents are essential for designing and testing systems that interact with humans, such as autonomous vehicles (AVs). These agents serve various purposes, from benchmarking AV performance to stress-testing the system’s limits, but all use cases share a key requirement: reliability. A simulation agent should behave as intended by the designer, minimizing unintended actions like collisions that can compromise the signal-to-noise ratio of analyses. As a foundation for reliable sim agents, we propose scaling self-play to thousands of scenarios on the Waymo Open Motion Dataset under semi-realistic limits on human perception and control. Training from scratch on a single GPU, our agents nearly solve the full training set within a day. They generalize effectively to unseen test scenes, achieving a 99.8% goal completion rate with less than 0.8% combined collision and off-road incidents across 10,000 held-out scenarios. Beyond in-distribution generalization, our agents show partial robustness to out-of-distribution scenes and can be fine-tuned in minutes to reach near-perfect performance in those cases. Demonstrations of agent behaviors can be found at this link. We open-source both the pre-trained agents and the complete code base. Demonstrations of agent behaviors can be found at https://sites.google.com/view/reliable-sim-agents.
@@ 2.5 Reward function
Here, $\mathbb{I}[.]$ is an indicator function that equals 1 if the condition is true and 0 otherwise. We assign weights $w_{\text{Goal Achieved}}=1.0$, $w_{\text{Collided}}=0.5$, and $w_{\text{Offroad}}=0.5$. Under our discount factor of .99.99.99.99, an agent will only be willing to accept a collision or off-road event if it would allow them to achieve their goal over $70$ steps faster. Note that such scenarios do occur in the dataset infrequently and this particular combination of collision weights and discount factor may occasionally make colliding an optimal behavior. An agent achieves the goal position when it reaches within a 2-meter radius of the target $(x,y)$. Once an agent reaches its goal, we remove it from the scene. This latter choice is made as it is ill-defined what an agent should do after it reaches its goal.
@@ 2.6 Collision behavior
During training and testing, we allow agents to continue the episode even after going off-road or colliding with another agent in the scene. Agents receive a penalty for each collision or off-road event, allowing them to accrue multiple penalties throughout an episode. A detailed discussion on can be found in Appendix B.1.
@@ 4.1 Scaling with data
We investigate whether agents with a partial view of the environment can solve all scenarios in the Waymo Open Motion Dataset. Our results show that nearly all scenarios can be solved successfully. After $1$ billion training steps, agents achieve a goal-reaching rate of $99.84$%, a collision rate of $0.38$%, and an off-road rate of $0.26$% on the training dataset. Furthermore, as shown in Figure 5, we observe that the failure rates would continue to tick down with continued training. This training run took $24$ hours on a single NVIDIA A100 GPU. [^3] Our metrics are computed in alignment with the way they are defined in WOSAC, but it should be noted that this is an over-optimistic metric as it includes many agents that simply need to remain in place as they are initialized right next to their goals. Excluding such agents, our metrics are: $99.20$% goal-reaching rate, $1.8$% collision rate, and $0.6$% offroad rate.
The agent-based metrics are similar to the scene-based metrics reported above: a goal rate of $99.72$%, collision rate of $0.26$%, and an off-road rate of $0.35\%$.
Figure 5: Batch performance throughout training. Left: Total number of collision and off-road events per batch during training. After 2 billion steps, the number of collisions stabilizes at approximately 15, which corresponds to 0.4% of the 4000 controlled agents in a batch. Center: Goal achievement rate per batch during training. This figure reports scene-level metrics, excluding mostly static agents counted in WOSAC that make the values appear higher. Right: The simulator is updated with new scenarios every 2 million steps in batches of 800 (uniform sampling with replacement). Over time, all scenarios are seen at least once.
@@
To assess how self-play performance scales with the diversity of training scenes, we conduct experiments with 100, 1,000, 10,000, and 100,000 unique training scenarios. Table 1 summarizes the results. We find no significant train-test gap when training with 10,000 scenarios or more, indicating the model generalizes well to new, unseen situations. Figure 4 shows the key metrics as a function of training dataset size. Notably, with 10,000 training scenarios, the model reaches nearly the ceiling of our benchmark, achieving a 99.81% goal-reaching rate, 0.44% collision rate, and 0.31% off-road rate on 10,000 held-out test scenarios.
@@ 4.2 Distribution of errors and remaining failure modes
We analyze scenarios that are not perfectly solved, defined as those with a collision rate or off-road rate greater than 0, or where at least one agent fails to reach its goal. A selection of failure modes can be viewed on the project page. Together, these account for $8.95$% of the test dataset ($896$ out of $10,000$ scenarios). Figure 6 shows the histogram of error distributions, revealing that most unsolved scenarios have only a small error rate. To examine potential relationships between failure modes, we compute the Pearson correlation between off-road fractions and collision rates. The result, $\rho=0.0135$, is not significant at $\alpha=0.05$, indicating no meaningful correlation between these two metrics in the unsolved scenarios and suggesting that errors are spread across scenarios.
@@ 4.3.1 Navigating backwards
Beyond generalization to within distribution scenarios, as reported in Section 4.1, we are interested in agent performance in out-of-distribution events. This is useful to know as researchers may typically manipulate scenarios or make them harder in some way to test the limits of AV systems. Where do these agents break, and how easily can they be finetuned? Driving backward, or navigating to goals behind agents is one such behavior that is rarely observed in the data. To quantify this, we analyzed the full training dataset ($\approx 129,000$ scenes) or about $4.2$ million controllable agents. Of these, we found approximately $30,000$ agents ($0.73$%) making a U-turn, and $47,000$ agents ($1.13$%) driving in reverse (see Appendix C.1 for the exact definition of these events). Further, most agents driving in reverse were simply pulling out of park, with goals immediately behind them, We observed a distinct lack of goals where the agent needs to execute a complex U-turn, making it plausibly out of distribution. We then hand-designed 13 scenarios from the test dataset with a total of 27 agents across all scenes, placing goals behind agents. This was done by setting the new goal for each agent to $(x_{f}-x_{i},y_{f}-y_{i})$, where $(x_{i},y_{i})$ is the initial position, and $(x_{f},y_{f})$ is the original goal. We chose the scenes in such a way that doing this process for all controlled agents results in valid and reachable goals. Figure 7.2 illustrates an example of such a scene.
@@ 1. Are these agents reliable enough?
Our agents are trained to optimize performance over given criteria above maximizing human likeness, making it unclear how closely they resemble real road users. A key direction for future work is balancing reliability with realism, ensuring agents not only meet performance standards but also accurately reflect human driving behavior across diverse scenarios.
@@ 6 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of the work, none which we feel must be specifically highlighted here.
@@ Appendix A Observation features and design choices
Appendix B Considerations for learning sim agents through self-play PPO
B.1 Collision behavior
@@
Appendix C Analyses.
C.1 Detecting out of distribution events
@@
C.2 Effect of initialization modes
report different initialization modes and trivial agents
Appendix D PPO implementation details.
D.1 Hyperparameters
@@
batch_size 262,144 Number of timesteps collected in each rollout.
@@
collision_weight -0.5 Penalty weight for collision events.
off_road_weight -0.5 Penalty weight for off-road events.
@@
Appendix E Compute resources

### v2

@@ Abstract
Simulation agents are essential for designing and testing systems that interact with humans, such as autonomous vehicles (AVs). These agents serve various purposes, from benchmarking AV performance to stress-testing system limits, but all applications share one key requirement: reliability. To enable systematic experimentation, a simulation agent must behave as intended. It should minimize actions that may lead to undesired outcomes, such as collisions, which can distort the signal-to-noise ratio in analyses. As a foundation for reliable sim agents, we propose scaling self-play to thousands of scenarios on the Waymo Open Motion Dataset under semi-realistic limits on human perception and control. Training from scratch on a single GPU, our agents nearly solve the full training set within a day. They generalize effectively to unseen test scenes, achieving a 99.8% goal completion rate with less than 0.8% combined collision and off-road incidents across 10,000 held-out scenarios. Beyond in-distribution generalization, our agents show partial robustness to out-of-distribution scenes and can be fine-tuned in minutes to reach near-perfect performance in those cases. We open-source the pre-trained agents and integrate them with a batched multi-agent simulator. Demonstrations of agent behaviors can be found at https://sites.google.com/view/reliable-sim-agents.
@@ 2.5 Reward function
Here, $\mathbb{I}[.]$ is an indicator function that equals 1 if the condition is true and 0 otherwise. We assign weights $w_{\text{Goal Achieved}}=1.0$, $w_{\text{Collided}}=0.75$, and $w_{\text{Offroad}}=0.75$. An agent achieves the goal position when it reaches within a 2-meter radius of the target $(x,y)$. Once an agent reaches its goal, we remove it from the scene. This latter choice is made as it is ill-defined what an agent should do after it reaches its goal.
@@ 2.6 Collision behavior
During training and testing, we allow agents to continue the episode even after going off-road or colliding with another agent in the scene. Agents receive a penalty for each collision or off-road event, allowing them to accrue multiple penalties throughout an episode. A detailed discussion on can be found in Appendix C.1.
@@ 4.1 Scaling with data
We investigate whether agents with a partial view of the environment can solve all scenarios in the Waymo Open Motion Dataset. Our results show that nearly all scenarios can be solved successfully. After $1$ billion training steps, agents achieve a goal-reaching rate of $99.84$%, a collision rate of $0.38$%, and an off-road rate of $0.26$% on the training dataset. Furthermore, as depicted in Figure 5, zooming in on the final four hours of training suggests that metrics exhibit a continued, albeit gradual, improvement, indicating that performance can be further increased with additional training. This training run took $24$ hours on a single NVIDIA A100 GPU. [^3] These metrics are computed in alignment with the way they are defined in WOSAC, but it should be noted that this is an over-optimistic metric as it includes many agents that simply need to remain in place as they are initialized right next to their goals. This initialization mode can be reproduced in the simulator by setting init_mode = all_objects. Excluding such agents, the performance metrics are: $99.40$% goal-reaching rate, $0.5$% collision rate, and $0.6$% offroad rate. The latter initialization mode, referred to as init_mode = all_non_trivial, only controls agents that must drive more than 2 meters before reaching their goal and is used during training.
The agent-based metrics are similar to the scene-based metrics reported above: a goal rate of $99.72$%, a collision rate of $0.26$%, and an off-road rate of $0.35\%$. Sample rollouts with the best-trained policy are shown in Figures 8, 9 and on the project page.
Figure 5: Batch performance throughout training. Left: Average reward per agent (maximum of 1) as a function of wall-clock time. We train agents for at most 24 hours. Center: Goal achievement rate per batch as a function of global steps (2 billion steps generated in 24 hours). Right: Percentage of agents that collide with another agent (red) or with a road edge (orange). All curves are smoothed using a rolling window of 250 steps. The inset figures show a zoomed-in view of the final four hours of the run, with the y-axes displayed on a logarithmic scale. The red annotations on the insets indicate the minimum and maximum values within the zoomed-in window. Note that the metrics reported during training are by excluding trivial agents, we only control agents that have to drive for more than 2 meters to reach their goal destination.
@@
We conduct experiments with 100, 1,000, 10,000, and 100,000 unique training scenarios to assess how self-play performance scales with the diversity of training scenes. Table 1 summarizes the results. We find no significant train-test gap when training with 10,000 scenarios or more, indicating the model generalizes well to new, unseen situations. Figure 4 shows the key metrics as a function of training dataset size. Notably, with 10,000 training scenarios, the model reaches nearly the ceiling of our benchmark, achieving a 99.81% goal-reaching rate, 0.44% collision rate, and 0.31% off-road rate on 10,000 held-out test scenarios.
@@ 4.2 Distribution of errors and remaining failure modes
We analyze scenarios that are not perfectly solved, defined as those with a collision rate or off-road rate greater than 0, or where at least one agent fails to reach its goal. A selection of failure modes can be viewed on the project page. Together, these account for $8.95$% of the test dataset ($896$ out of $10,000$ scenarios). Figure 6 shows the histogram of error distributions, revealing that most unsolved scenarios have only a small error rate. We compute the Pearson correlation between off-road fractions and collision rates to examine potential relationships between failure modes. The result, $\rho=0.0135$, is not significant at $\alpha=0.05$, indicating no meaningful correlation between these two metrics in the unsolved scenarios and suggesting that errors are spread across scenarios.
@@ 4.3.1 Navigating backwards
Beyond generalization to within distribution scenarios, as reported in Section 4.1, we are interested in agent performance in out-of-distribution events. This is useful to know as researchers may typically manipulate scenarios or make them harder in some way to test the limits of AV systems. Where do these agents break, and how easily can they be finetuned? Driving backward, or navigating to goals behind agents is one such behavior that is rarely observed in the data. To quantify this, we analyzed the full training dataset ($\approx 129,000$ scenes) or about $4.2$ million controllable agents. Of these, we found approximately $30,000$ agents ($0.73$%) making a U-turn, and $47,000$ agents ($1.13$%) driving in reverse (see Appendix D.1 for the exact definition of these events). Further, most agents driving in reverse were simply pulling out of park, with goals immediately behind them, We observed a distinct lack of goals where the agent needs to execute a complex U-turn, making it plausibly out of distribution. We then hand-designed 13 scenarios from the test dataset with a total of 27 agents across all scenes, placing goals behind agents. This was done by setting the new goal for each agent to $(x_{f}-x_{i},y_{f}-y_{i})$, where $(x_{i},y_{i})$ is the initial position, and $(x_{f},y_{f})$ is the original goal. We chose the scenes in such a way that doing this process for all controlled agents results in valid and reachable goals. Figure 7.2 illustrates an example of such a scene.
@@ 1. Are these agents reliable enough?
Our agents are trained to optimize performance over given criteria above maximizing human likeness, making it unclear how closely they resemble real road users. An interesting direction for future work is balancing reliability with realism, ensuring agents not only meet performance standards but also accurately reflect human driving behavior across diverse scenarios.
5.2 Concluding thoughts
@@
We demonstrated the potential of scaling self-play to develop agents that can be precisely controlled to meet specific criteria in autonomous driving. While not explored in this paper, we anticipate that our findings extend to other domains such as neuroscience, where agent-based modeling is gaining momentum (Aldarondo et al., 2024; Johnson-Yu et al., ; Castro et al., 2025). In neuroscience, researchers are increasingly using physics-based simulators to create digital twins of animals, enabling cost-effective and controlled experimentation. For these agents to be useful models of animal behavior, reliability and robustness appear essential. A rodent foraging model, for example, should not exhibit free movement. We hope our work contributes to the improvement of agent-based modeling, helping to enhance controllability and robustness across different domains.
@@ 6 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of the work, none of which we feel must be specifically highlighted here.
@@ Appendix A Observation features and design choices
Appendix B Additional figures
B.1 Sample rollouts
Figure 8: Example rollouts with the best-trained policy. Agents controlled by the trained policy are shown in blue, while static agents are colored in grey.
Figure 9: Example rollouts with the best-trained policy. Agents controlled by the trained policy are shown in blue, while static agents are colored in grey.
Appendix C Considerations for learning sim agents through self-play PPO
C.1 Collision behavior
@@
Appendix D Analyses.
D.1 Detecting out of distribution events
@@
Appendix E PPO implementation details.
E.1 Hyperparameters
@@
batch_size 524,288 Number of timesteps collected in each rollout.
@@
collision_weight -0.75 Penalty weight for collision events.
off_road_weight -0.75 Penalty weight for off-road events.
@@
Appendix F Compute resources